# Multi-Agent Dynamic Relational Reasoning for Social Robot Navigation

### v1

@@ Abstract
Social robot navigation can be helpful in various contexts of daily life but requires safe human-robot interactions and efficient trajectory planning. While modeling pairwise relations has been widely studied in multi-agent interacting systems, the ability to capture larger-scale group-wise activities is limited. In this paper, we propose a systematic relational reasoning approach with explicit inference of the underlying dynamically evolving relational structures, and we demonstrate its effectiveness for multi-agent trajectory prediction and social robot navigation. In addition to the edges between pairs of nodes (i.e., agents), we propose to infer hyperedges that adaptively connect multiple nodes to enable group-wise reasoning in an unsupervised manner. Our approach infers dynamically evolving relation graphs and hypergraphs to capture the evolution of relations, which the trajectory predictor employs to generate future states. Meanwhile, we propose to regularize the sharpness and sparsity of the learned relations and the smoothness of the relation evolution, which proves to enhance training stability and model performance. The proposed approach is validated on synthetic crowd simulations and real-world benchmark datasets. Experiments demonstrate that the approach infers reasonable relations and achieves state-of-the-art prediction performance. In addition, we present a deep reinforcement learning (DRL) framework for social robot navigation, which incorporates relational reasoning and trajectory prediction systematically. In a group-based crowd simulation, our method outperforms the strongest baseline by a significant margin in terms of safety, efficiency, and social compliance in dense, interactive scenarios.
@@
We validate the trajectory prediction method with relational reasoning on both crowd simulations and real-world benchmark datasets. Our method infers reasonable evolving hypergraphs for group-wise relational reasoning and achieves state-of-the-art performance. We also evaluate the DRL-based robot navigation method on our group-based crowd simulations, which demonstrates enhanced navigation performance, social awareness and appropriateness in robot motion planning.
@@ II-B Social Robot Navigation
Social robot navigation not only considers human behavior modeling but also robot motion planning [36]. The robot navigation system should consider the safety and efficiency of the robot and humans in the environment. Besides developing effective human behavior modeling methods for trajectory prediction to capture the interactions, another challenge is how to ensure the safety of humans during navigation. On one hand, some work focuses on game-theoretic approaches. For example, the human-robot interaction problem can be formulated as a non-cooperative non-zero-sum game [37]. Multi-agent interaction problems can be formulated as a repeated linear-quadratic game [38]. On the other hand, some work proposes reinforcement learning based methods to implicitly capture the human-robot interactions and synthesize the robot planning policy directly based on observations [39, 40, 41, 42, 13]. Our method falls into the second category, which incorporates interaction modeling and human behavior modeling into the reinforcement learning framework systematically.
@@ V Social Robot Navigation
In this section, we adopt the same notations for trajectories from Section IV. Given the historical positions of agent (i.e., ), we employ our method in Section IV to predict the future positions of the agent as . We denote the robot’s states as , which includes current robot position , the goal position , the robot’s velocity , the maximum robot speed , and the robot radius . We define the state of the MDP as
@@ V-A Overview
Fig. 4 illustrates the overall pipeline of our robot navigation method, which integrates deep reinforcement learning with multi-agent relational reasoning and trajectory prediction. The pipeline consists of our trajectory prediction model, two attention modules (i.e., human-human attention and robot-human attention) for feature aggregation adapted from CrowdNav++ [13], and a recurrent layer to output value and policy. Specifically, the trajectory predictor processes the historical trajectories of observable agents to generate their future state hypotheses. These predictions are first transformed by a human-human attention module into a human-centric embedding. Then, these embeddings, along with the current robot state, are processed by the robot-human attention module, resulting in a refined feature representation of human agents considering human-robot interactions. This enhanced representation, concatenated with the current robot state, is then fed into a recurrent layer, which outputs the policy and the value .
@@ V-B Attention Modules
Attention layers in our model play an important role by assigning weights to each edge connecting to an agent, thereby emphasizing important interactions. The human-human and robot-human attention layers employ the scaled dot-product attention technique.
In the human-human attention layer, the current and predicted future states of humans are concatenated and passed through linear layers to derive the query, key, and value:
where represents the concatenated state for all agents. The attention weights are obtained from , and the human embeddings at the current time step are derived using a multi-head dot-product attention mechanism with eight attention heads.
In the robot-human attention layer, the human embeddings obtained from the human-human attention layer serve as inputs for deriving the query and value:
@@ V-D Reward Function
Collision and prediction rewards. To encourage safe navigation, we penalize collisions and potential intrusions into the space of humans. The prediction reward penalizes the robot based on its proximity to the predicted future positions of humans, which is written as
@@
Group interaction reward. To minimize discomfort to individuals and discourage intrusion into human groups, we consider the geometry of group formations. Fig. 5 illustrates the scenarios where the robot complies with or violates the social norms. For each group, we calculate a convex hull encompassing all agents within the group. The reward is defined as a function of the robot’s distance to the closest point on the group’s convex hull, normalized by the area of the convex hull, which is written as
where is a scaling factor, and represents the convex hull of a group.
@@ VI-A1 Human Trajectory Datasets
ETH-UCY [63, 64]: These datasets were collected in real-world traffic scenarios where pedestrians move on the street or in the university campus. Some scenarios are crowded and full of pairwise and group-wise interactions between pedestrians who naturally form groups and move together. We used the standard split of training, validation, and testing datasets. Moreover, we also selected the test cases with more than eight pedestrians to construct a difficult test subset to compare the model performance in dense scenarios with strong interactions.
@@ VI-A2 Group-Aware Crowd Simulation
We evaluate social robot navigation with a crowd simulation as an interactive environment for reinforcement learning, which is also used to collect a synthetic dataset. We develop a group-based crowd simulator that incorporates the Social Force model [18] to enable the simulation of crowd dynamics with human groups. More specifically, we design a specific scenario to observe changes in group behaviors. The simulation environment is a 2D square space (), where we randomly initialize five groups of pedestrians. Each group, containing up to five individuals, starts within the bounded area. This setup allows for a maximum of 25 pedestrians, leading to dense and interactive crowd scenarios. Every group is assigned a random destination along the neighboring or the opposite boundary and moves at a random maximum speed () ranging from to . Each pedestrian is modeled with a radius () between and .
@@ VI-B1 Trajectory Prediction
minADE: the minimum distance among trajectory samples between the predicted trajectories and the ground truth over all the entities in the prediction horizon.
minFDE: the minimum deviated distance among trajectory samples at the last predicted time step.
@@ VI-C2 Social Robot Navigation
Dataset |
MID [14] |
GroupNet [12] |
Graph-TERN [15] |
EigenTraj [16] |
SCG |
SHG |
SCG + SHG |
DCG + DHG |
DCG + DHG + SM |
DCG + DHG + SM + SH |
DCG + DHG + SM + SH + SP (Equal Attention) |
DCG + DHG + SM + SH + SP |
ETH |
0.39 / 0.66 |
0.46 / 0.73 |
0.42 / 0.58 |
0.36 / 0.57 |
0.41 / 0.68 |
0.45 / 0.76 |
0.40 / 0.66 |
0.37 / 0.63 |
0.35 / 0.62 |
0.33 / 0.59 |
0.34 / 0.60 |
0.33 / 0.54 |
HOTEL |
0.13 / 0.22 |
0.15 / 0.25 |
0.14 / 0.23 |
0.13 / 0.21 |
0.16 / 0.27 |
0.19 / 0.28 |
0.16 / 0.24 |
0.15 / 0.23 |
0.13 / 0.21 |
0.13 / 0.18 |
0.14 / 0.21 |
0.12 / 0.17 |
UNIV |
0.22 / 0.45 |
0.26 / 0.49 |
0.26 / 0.45 |
0.24 / 0.43 |
0.24 / 0.49 |
0.28 / 0.51 |
0.22 / 0.47 |
0.21 / 0.45 |
0.20 / 0.43 |
0.18 / 0.41 |
0.19 / 0.40 |
0.18 / 0.36 |
ZARA1 |
0.17 / 0.30 |
0.21 / 0.39 |
0.21 / 0.37 |
0.20 / 0.35 |
0.23 / 0.36 |
0.25 / 0.40 |
0.22 / 0.35 |
0.21 / 0.32 |
0.20 / 0.31 |
0.18 / 0.30 |
0.19 / 0.29 |
0.17 / 0.29 |
ZARA2 |
0.13 / 0.27 |
0.17 / 0.33 |
0.17 / 0.29 |
0.14 / 0.25 |
0.21 / 0.30 |
0.23 / 0.31 |
0.20 / 0.27 |
0.18 / 0.24 |
0.17 / 0.20 |
0.15 / 0.21 |
0.16 / 0.23 |
0.13 / 0.20 |
AVG |
0.21 / 0.38 |
0.25 / 0.44 |
0.24 / 0.38 |
0.21 / 0.36 |
0.24 / 0.45 |
0.26 / 0.47 |
0.23 / 0.42 |
0.21 / 0.39 |
0.20 / 0.35 |
0.19 / 0.32 |
0.21 / 0.36 |
0.18 / 0.30 |
@@
Dataset |
MID [14] |
GroupNet [12] |
Graph-TERN [15] |
EigenTraj [16] |
SCG |
SHG |
SCG + SHG |
DCG + DHG |
DCG + DHG + SM |
DCG + DHG + SM + SH |
DCG + DHG + SM + SH + SP (Equal Attention) |
DCG + DHG + SM + SH + SP |
ETH |
0.42 / 0.77 |
0.49 / 0.89 |
0.52 / 0.84 |
0.40 / 0.73 |
0.45 / 0.79 |
0.54 / 0.96 |
0.43 / 0.72 |
0.41 / 0.71 |
0.38 / 0.65 |
0.36 / 0.63 |
0.36 / 0.65 |
0.35 / 0.61 |
HOTEL |
0.15 / 0.25 |
0.18 / 0.30 |
0.17 / 0.28 |
0.14 / 0.24 |
0.20 / 0.31 |
0.24 / 0.35 |
0.17 / 0.29 |
0.17 / 0.26 |
0.15 / 0.24 |
0.14 / 0.23 |
0.15 / 0.25 |
0.14 / 0.20 |
UNIV |
0.23 / 0.55 |
0.27 / 0.58 |
0.29 / 0.61 |
0.26 / 0.54 |
0.28 / 0.55 |
0.33 / 0.69 |
0.26 / 0.54 |
0.24 / 0.51 |
0.23 / 0.48 |
0.22 / 0.46 |
0.22 / 0.49 |
0.21 / 0.45 |
ZARA1 |
0.21 / 0.38 |
0.26 / 0.45 |
0.23 / 0.42 |
0.22 / 0.37 |
0.25 / 0.41 |
0.29 / 0.52 |
0.23 / 0.38 |
0.23 / 0.36 |
0.21 / 0.35 |
0.20 / 0.33 |
0.21 / 0.33 |
0.19 / 0.31 |
ZARA2 |
0.20 / 0.34 |
0.21 / 0.36 |
0.21 / 0.32 |
0.18 / 0.28 |
0.26 / 0.32 |
0.28 / 0.41 |
0.23 / 0.29 |
0.21 / 0.28 |
0.19 / 0.25 |
0.18 / 0.24 |
0.19 / 0.26 |
0.16 / 0.23 |
AVG |
0.31 / 0.50 |
0.29 / 0.49 |
0.27 / 0.47 |
0.25 / 0.44 |
0.30 / 0.56 |
0.34 / 0.65 |
0.28 / 0.50 |
0.25 / 0.45 |
0.23 / 0.41 |
0.22 / 0.39 |
0.24 / 0.43 |
0.20 / 0.34 |
@@
Dataset |
MID [14] |
GroupNet [12] |
Graph-TERN [15] |
EigenTraj [16] |
SCG |
SHG |
SCG + SHG |
DCG + DHG |
DCG + DHG + SM |
DCG + DHG + SM + SH |
DCG + DHG + SM + SH + SP (Equal Attention) |
DCG + DHG + SM + SH + SP |
SDD |
9.7 / 15.3 |
9.7 / 15.3 |
8.4 / 14.3 |
8.1 / 13.1 |
12.7 / 22.5 |
14.0 / 25.8 |
12.0 / 20.4 |
9.5 / 15.4 |
8.8 / 14.7 |
8.4 / 13.6 |
9.4 / 14.9 |
8.1 / 12.8 |
NBA |
1.65 / 2.66 |
1.53 / 2.45 |
1.46 / 2.19 |
1.49 / 2.30 |
1.69 / 2.73 |
1.80 / 3.04 |
1.56 / 2.32 |
1.37 / 2.00 |
1.28 / 1.81 |
1.23 / 1.73 |
1.31 / 1.87 |
1.16 / 1.64 |

### v2

@@ Abstract
Social robot navigation can be helpful in various contexts of daily life but requires safe human-robot interactions and efficient trajectory planning. While modeling pairwise relations has been widely studied in multi-agent interacting systems, the ability to capture larger-scale group-wise activities is limited. In this paper, we propose a systematic relational reasoning approach with explicit inference of the underlying dynamically evolving relational structures, and we demonstrate its effectiveness for multi-agent trajectory prediction and social robot navigation. In addition to the edges between pairs of nodes (i.e., agents), we propose to infer hyperedges that adaptively connect multiple nodes to enable group-wise reasoning in an unsupervised manner. Our approach infers dynamically evolving relation graphs and hypergraphs to capture the evolution of relations, which the trajectory predictor employs to generate future states. Meanwhile, we propose to regularize the sharpness and sparsity of the learned relations and the smoothness of the relation evolution, which proves to enhance training stability and model performance. The proposed approach is validated on synthetic crowd simulations and real-world benchmark datasets. Experiments demonstrate that the approach infers reasonable relations and achieves state-of-the-art prediction performance. In addition, we present a deep reinforcement learning (DRL) framework for social robot navigation, which incorporates relational reasoning and trajectory prediction systematically. In a group-based crowd simulation, our method outperforms the strongest baseline by a significant margin in terms of safety, efficiency, and social compliance in dense, interactive scenarios. We also demonstrate the practical applicability of our method with real-world robot experiments. The code and experiment videos can be found on the project website: https://relational-reasoning-nav.github.io/.
@@
We validate the trajectory prediction method with relational reasoning on both crowd simulations and real-world benchmark datasets. Our method infers reasonable evolving hypergraphs for group-wise relational reasoning and achieves state-of-the-art performance. We also evaluate the DRL-based robot navigation method on our group-based crowd simulations and real-world scenarios, demonstrating enhanced navigation performance, social awareness, and appropriateness in robot motion planning.
@@ II-B Social Robot Navigation
Social robot navigation not only considers human behavior modeling but also robot motion planning [36]. The robot navigation system should consider the safety and efficiency of the robot and humans in the environment. Besides developing effective human behavior modeling methods for trajectory prediction to capture the interactions, another challenge is how to ensure the safety of humans during navigation. On one hand, some work focuses on game-theoretic approaches. For example, the human-robot interaction problem can be formulated as a non-cooperative non-zero-sum game [37]. Multi-agent interaction problems can be formulated as a repeated linear-quadratic game [38]. On the other hand, some work proposes reinforcement learning based methods to implicitly capture the human-robot interactions and synthesize the robot planning policy directly based on the observations [39, 40, 41, 42, 13]. Liu et al. [13] propose an RL-based social navigation method with trajectory prediction of surrounding pedestrians. However, their prediction and planning models do not consider group-wise relations or activities. Katyal et al. [42] propose to learn a group-aware policy for robot navigation. However, their method only uses the historical state information for robot decision making while not considering the future behaviors of pedestrians. Also, their method requires ground truth group membership information during training and does not consider the evolution of group relations over time. In contrast, our approach explicitly infers dynamic pairwise and group-wise relations in an unsupervised manner, which incorporates interaction modeling and human behavior modeling into the reinforcement learning framework systematically.
@@ V Social Robot Navigation
In this section, we adopt the same notations for trajectories from Section IV. Given the historical positions of agent (i.e., ), we employ our method in Section IV to predict the future positions of the agent as . We denote the robot’s states as , which includes current robot position , the goal position , the robot’s velocity , the maximum robot speed , and the robot radius . Furthermore, to enhance the realism and complexity of the environment, we also account for static obstacles as a global state, denoted as . Each represents the state of static obstacle , which includes both its position and radius. In this work, we only consider circular static obstacles while more complex shapes are left for future work. We define the state of the MDP as
@@ V-A Overview
Fig. 4 illustrates the overall pipeline of our robot navigation method, which integrates deep reinforcement learning with multi-agent relational reasoning and trajectory prediction. The pipeline consists of our trajectory prediction model, two attention modules (i.e., human-human/obstacle attention and robot-human/obstacle attention) for feature aggregation adapted from CrowdNav++ [13], and a recurrent layer to output value and policy. Specifically, the trajectory predictor processes the historical trajectories of observable agents to generate their future state hypotheses. These predictions are first transformed by a human-human/obstacle attention module into a human-centric embedding. Then, these embeddings, along with the current robot state, are processed by the robot-human attention module, resulting in a refined feature representation of human agents considering human-robot interactions. This enhanced representation, concatenated with the current robot state, is then fed into a recurrent layer, which outputs the policy and the value .
@@ V-B Attention Modules
Attention layers in our model play an important role by assigning weights to each edge connecting to an agent, thereby emphasizing important interactions. The human-human/obstacle and robot-human/obstacle attention layers employ the scaled dot-product attention technique.
In the human-human/obstacle attention layer, the current and predicted future states of humans are concatenated and passed through linear layers to derive the query, key, and value:
where represents the concatenated state for all agents. The attention weights are obtained from , and the human embeddings at the current time step are derived using a multi-head dot-product attention mechanism with eight attention heads. Note that the attention weights between static obstacles are set to since they do not influence each other.
In the robot-human/obstacle attention layer, the human embeddings obtained from the human-human/obstacle attention layer serve as inputs for deriving the query and value:
@@ V-D Reward Function
Human collision and prediction rewards. To encourage safe navigation, we penalize collisions and potential intrusions into the space of humans. The prediction reward penalizes the robot based on its proximity to the predicted future positions of humans, which is written as
@@
Obstacle collision reward. Similar to the human collision and prediction rewards, we have a distance reward that penalizes the robot based on its proximity to the nearest obstacles, as expressed by:
where is a scaling factor. The collision reward is applied when the robot collides with any obstacle, imposing a heavy penalty for such events. We set lower penalty weights for and compared to those associated with humans to emphasize safe navigation around humans.
Group interaction reward. To minimize discomfort to individuals and discourage intrusion into human groups, we consider the geometry of group formations. Fig. 5 illustrates the scenarios where the robot complies with or violates the social norms. For each group, we calculate a convex hull encompassing all agents within the group. For example, with a group of pedestrian , where is the number of pedestrians in this group, we define the convex hull of this group as
i.e., the convex hull is the intersection of all convex sets that contain . We employ the Qhull algorithm [62] to construct these convex hulls by initially selecting the extreme points based on -coordinates and dividing the points into two subsets. Using a recursive divide-and-conquer approach, it identifies the farthest point from each line segment, refining the subsets and expanding the convex boundary until no external points remain. This method can efficiently delineate the smallest convex polygon enclosing all points.
The reward is defined as a function of the robot’s distance to the closest point on the group’s convex hull, normalized by the area of the convex hull, which is written as
here the is the shortest distance from the robot to the convex hull, is a scaling factor. is a condition variable, which is set to when the robot is located outside the convex hull. Conversely, it is set to to incentivize the robot to move out from the group with the shortest path.
@@
where are the states of the robot reaching the goal, are the states of robot collision with a human, and are the states of robot collision with a static obstacle.
@@ VI-A1 Human Trajectory Datasets
ETH-UCY [64, 65]: These datasets were collected in real-world traffic scenarios where pedestrians move on the street or the university campus. Some scenarios are crowded and full of pairwise and group-wise interactions between pedestrians who naturally form groups and move together. We used the standard split of training, validation, and testing datasets. Moreover, we also selected the test cases with more than eight pedestrians to construct a difficult test subset to compare the model performance in dense scenarios with strong interactions.
@@ VI-A2 Group-Aware Crowd Simulation
We evaluate social robot navigation with a crowd simulation as an interactive environment for reinforcement learning, which is also used to collect a synthetic dataset. We develop a group-based crowd simulator that incorporates the ORCA model [67] to enable the simulation of crowd dynamics with human groups. More specifically, we design a specific scenario to observe changes in group behaviors. The simulation environment is a 2D square space (), where we randomly initialize five groups of pedestrians. For both training and evaluation, each group contains 2-5 individuals randomly, i.e., in total 10-25 pedestrians. For each group, all members will be positioned within a circular area. Different groups do not have overlapping initial spaces. Every group is assigned a random destination along the neighboring or the opposite boundary and moves at a random maximum speed () ranging from to . Each pedestrian is modeled with a radius () between and . In the square, a random number of static obstacles, ranging from one to the number of groups, will be randomly positioned within a circular area. The center of this area coincides with the center of the square, and it has a radius of , which is an inscribed circle. Each static obstacle is modeled with a radius between and .
@@ VI-B1 Trajectory Prediction
minADE20: the minimum distance among trajectory samples between the predicted trajectories and the ground truth over all the entities in the prediction horizon.
minFDE20: the minimum deviated distance among trajectory samples at the last predicted time step.
@@ VI-C2 Social Robot Navigation
DSOA [68]: This is a closed-form solution for dynamic spatial navigation, which creates smooth motion fields to ensure the robot navigates safely within confined spaces and around both static and dynamic obstacles.
@@
Group-Aware [42]: This model employs deep reinforcement learning to develop navigation policies that recognize and respect dynamic human group formations, which enhances robot navigation by reducing collisions and adhering to social norms.
@@
Dataset | MID [14] | GroupNet [12] | Graph-TERN [15] | EigenTraj [16] | SCG | SHG | SCG + SHG | DCG + DHG | DCG + DHG + SM | DCG + DHG + SM + SH | DCG + DHG + SM + SH + SP (Equal Attention) | DCG + DHG + SM + SH + SP |
ETH | 0.39 / 0.66 | 0.46 / 0.73 | 0.42 / 0.58 | 0.36 / 0.57 | 0.41 / 0.68 | 0.45 / 0.76 | 0.40 / 0.66 | 0.37 / 0.63 | 0.35 / 0.62 | 0.33 / 0.59 | 0.34 / 0.60 | 0.33 / 0.54 |
HOTEL | 0.13 / 0.22 | 0.15 / 0.25 | 0.14 / 0.23 | 0.13 / 0.21 | 0.16 / 0.27 | 0.19 / 0.28 | 0.16 / 0.24 | 0.15 / 0.23 | 0.13 / 0.21 | 0.13 / 0.18 | 0.14 / 0.21 | 0.12 / 0.17 |
UNIV | 0.22 / 0.45 | 0.26 / 0.49 | 0.26 / 0.45 | 0.24 / 0.43 | 0.24 / 0.49 | 0.28 / 0.51 | 0.22 / 0.47 | 0.21 / 0.45 | 0.20 / 0.43 | 0.18 / 0.41 | 0.19 / 0.40 | 0.18 / 0.36 |
ZARA1 | 0.17 / 0.30 | 0.21 / 0.39 | 0.21 / 0.37 | 0.20 / 0.35 | 0.23 / 0.36 | 0.25 / 0.40 | 0.22 / 0.35 | 0.21 / 0.32 | 0.20 / 0.31 | 0.18 / 0.30 | 0.19 / 0.29 | 0.17 / 0.29 |
ZARA2 | 0.13 / 0.27 | 0.17 / 0.33 | 0.17 / 0.29 | 0.14 / 0.25 | 0.21 / 0.30 | 0.23 / 0.31 | 0.20 / 0.27 | 0.18 / 0.24 | 0.17 / 0.20 | 0.15 / 0.21 | 0.16 / 0.23 | 0.13 / 0.20 |
AVG | 0.21 / 0.38 | 0.25 / 0.44 | 0.24 / 0.38 | 0.21 / 0.36 | 0.24 / 0.45 | 0.26 / 0.47 | 0.23 / 0.42 | 0.21 / 0.39 | 0.20 / 0.35 | 0.19 / 0.32 | 0.21 / 0.36 | 0.18 / 0.30 |
@@
Dataset | MID [14] | GroupNet [12] | Graph-TERN [15] | EigenTraj [16] | SCG | SHG | SCG + SHG | DCG + DHG | DCG + DHG + SM | DCG + DHG + SM + SH | DCG + DHG + SM + SH + SP (Equal Attention) | DCG + DHG + SM + SH + SP |
ETH | 0.42 / 0.77 | 0.49 / 0.89 | 0.52 / 0.84 | 0.40 / 0.73 | 0.45 / 0.79 | 0.54 / 0.96 | 0.43 / 0.72 | 0.41 / 0.71 | 0.38 / 0.65 | 0.36 / 0.63 | 0.36 / 0.65 | 0.35 / 0.61 |
HOTEL | 0.15 / 0.25 | 0.18 / 0.30 | 0.17 / 0.28 | 0.14 / 0.24 | 0.20 / 0.31 | 0.24 / 0.35 | 0.17 / 0.29 | 0.17 / 0.26 | 0.15 / 0.24 | 0.14 / 0.23 | 0.15 / 0.25 | 0.14 / 0.20 |
UNIV | 0.23 / 0.55 | 0.27 / 0.58 | 0.29 / 0.61 | 0.26 / 0.54 | 0.28 / 0.55 | 0.33 / 0.69 | 0.26 / 0.54 | 0.24 / 0.51 | 0.23 / 0.48 | 0.22 / 0.46 | 0.22 / 0.49 | 0.21 / 0.45 |
ZARA1 | 0.21 / 0.38 | 0.26 / 0.45 | 0.23 / 0.42 | 0.22 / 0.37 | 0.25 / 0.41 | 0.29 / 0.52 | 0.23 / 0.38 | 0.23 / 0.36 | 0.21 / 0.35 | 0.20 / 0.33 | 0.21 / 0.33 | 0.19 / 0.31 |
ZARA2 | 0.20 / 0.34 | 0.21 / 0.36 | 0.21 / 0.32 | 0.18 / 0.28 | 0.26 / 0.32 | 0.28 / 0.41 | 0.23 / 0.29 | 0.21 / 0.28 | 0.19 / 0.25 | 0.18 / 0.24 | 0.19 / 0.26 | 0.16 / 0.23 |
AVG | 0.31 / 0.50 | 0.29 / 0.49 | 0.27 / 0.47 | 0.25 / 0.44 | 0.30 / 0.56 | 0.34 / 0.65 | 0.28 / 0.50 | 0.25 / 0.45 | 0.23 / 0.41 | 0.22 / 0.39 | 0.24 / 0.43 | 0.20 / 0.34 |
@@
Dataset | MID [14] | GroupNet [12] | Graph-TERN [15] | EigenTraj [16] | SCG | SHG | SCG + SHG | DCG + DHG | DCG + DHG + SM | DCG + DHG + SM + SH | DCG + DHG + SM + SH + SP (Equal Attention) | DCG + DHG + SM + SH + SP |
SDD | 9.7 / 15.3 | 9.7 / 15.3 | 8.4 / 14.3 | 8.1 / 13.1 | 12.7 / 22.5 | 14.0 / 25.8 | 12.0 / 20.4 | 9.5 / 15.4 | 8.8 / 14.7 | 8.4 / 13.6 | 9.4 / 14.9 | 8.1 / 12.8 |
NBA | 1.65 / 2.66 | 1.53 / 2.45 | 1.46 / 2.19 | 1.49 / 2.30 | 1.69 / 2.73 | 1.80 / 3.04 | 1.56 / 2.32 | 1.37 / 2.00 | 1.28 / 1.81 | 1.23 / 1.73 | 1.31 / 1.87 | 1.16 / 1.64 |